OpenStudy (anonymous):

A stick of gum is 2.6 cm long a) How long is it if it stretched to 1.7 cm time its intial length? b) It's stretched to 2.3cm time its intial length before it breaks .How long is it then?

OpenStudy (ash2326):

iniitial length a-2.6 a) stretched to 1.7 times so length=-1.7*2.6=4.42 cm b) 2.3 times so 2.3*2.6=5.98 cm

OpenStudy (anonymous):

tk